OpenStudy (anonymous):

show that the equation x^2+y^2+x+k(x^2+y^2-2x+y-1)=0 represents a circle, and find the center and the radius

OpenStudy (anonymous):

what i did so far was.. \[x^2+y^2+x+k(x^2+y^2-2x+y-1)=0\]\[x^2+y^2+x+kx^2+ky^2-2kx+ky-k=0\]\[x^2+kx^2+y^2+ky^2+x-2kx+ky-k=0\]\[(1+k)x^2+(1+k)y^2+(1-2k)x+ky-k=0\]thats wat i did but im stuck there...

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

I think complete the square. Rearrange a little first \[[(1+k)x^2+(1-2k)x]+[(1+k)y^2+ky]=k\] divide by (1 + k) \[(x^2+x)+(y^2+\frac{ k }{ 1+k })=\frac{ k }{ 1+k }\]

OpenStudy (anonymous):

\[(x+\frac{ 1 }{ 2 })^2+\left( y+\frac{ k }{ 2(1+k) } \right)^2=\frac{ k }{ 1+k }+\frac{ 1 }{ 4 }+\frac{ k^2 }{ 4(1+k)^2 }\]

OpenStudy (anonymous):

that's really nasty, but once you combine the stuff on the right, the square root of it will be the radius

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] (h, k) = center and r = radius

OpenStudy (anonymous):

shdnt it be \[x^2+\frac{ 1-2k }{ 1+k }x+y^2+\frac{ k }{ 1+k }y=\frac{ k }{ 1+k }\]

OpenStudy (anonymous):

@peachpi

OpenStudy (campbell_st):

that's correct so it becomes \[(x + \frac{1 - 2k}{2(1 + k)})^2 + (y + \frac{k}{2(1 + k)})^2 = \frac{k}{1 + k} + (\frac{1 - 2k}{2(1 + k)})^2 + (\frac{k}{2(1 + k})^2 \] then just simplify the right hand side

OpenStudy (anonymous):

oh yeah, sorry. good catch

OpenStudy (anonymous):

thanks guys